OpenStudy (anonymous):

A triangle has a side with length 6 feet and another side with length 8 feet. The angle between the sides measures 73 degrees. Find the area of the triangle. Round your answer to the nearest tenth.

OpenStudy (raden):

area = 1/2 * 6 * 8 * sin73 = ...

OpenStudy (ddcamp):

 \[A = \frac{ 1 }{ 2 } A B \sin(c°)\]

OpenStudy (anonymous):

I got 23.04 squared feet as my answer. Did I do it right? Thank you both by the way! I wish I could give you both medals haha

OpenStudy (raden):

yep, that's right

OpenStudy (anonymous):

Thank you!

OpenStudy (raden):

yw